### v1

# A structure-preserving, second-order-in-time scheme for the von Neumann equation with power nonlinearity

Abstract.
In this paper we propose a structure-preserving, linearly implicit, second-order-in-time scheme for the numerical solution of the von Neumann equation with power nonlinearity (also known as the Alber equation). Fourth order finite differences are used for the spatial discretization. We highlight the importance of the correct initialization of the method in achieving the expected order of convergence in space and time. As illustrative examples, we investigate the bifurcation from Landau damping to modulation instability. In that context, amplification factors in the fully developed modulation instability for this nonlinear equation are computed for the first time.
Key words and phrases:
relaxation scheme, structure preserving, finite differences, Alber equation, modulation instability, nonlinear Landau damping.1. Introduction
The von Neumann equation with power nonlinearity and homogeneous background, also known as the Alber equation, is given by
where and are real-valued parameters, is a given autocorrelation function, and is Hermitian, The autocorrelation describes the known homogeneous background of the problem, and the unknown function controls the inhomogeneity.
Equation (1.1) appears in the context of ocean waves [1, 4, 9, 10, 12, 13, 14, 15, 16], and is known to exhibit a bifurcation between a stable regime (nonlinear Landau damping) and an unstable regime (modulation instability), depending on the homogeneous background In this paper, we introduce a second-order in time relaxation scheme, which inherits the problem’s key balance law at the discrete level. An advanced initialization for the scheme is introduced and validated against exact solutions. This is required to ensure that both the main and relaxation variable achieve the expected order of convergence. Finally, the scheme is used to investigate both modulation instability and nonlinear Landau damping for equation (1.1).
The nonlinearity of the equation (1.1) is considered very singular compared to other equations of similar structure, and is well known for presenting challenges in its analysis and numerical solution. Well-posedness and smoothness of solutions for the Alber equation (1.1) was rigorously established for the first time in [2].
1.1. Numerical methods for the Alber equation: state of the art
To the best of our knowledge, there are only two papers dealing with the numerical solution of the Alber equation. These are [14, 16], and both treat equation (1.1) after a change of variables, to “mean position” and “lag” (this change of variables first appears in [1], and it is employed in the calculations of Appendix A). Both papers [14, 16] use a forward Euler solver for time-stepping, and finite differences in space. Moreover, a non-standard boundary condition needs to be introduced for the variable
The works [14, 16] are fundamental for the field, and they highlight the challenges this non-standard equation poses. The introduction of novel boundary conditions in introduces one more challenging element in the problem and numerical method. Our approach of solving directly (1.1) with periodic boundary conditions in the variables preserves by construction the original physical context [11]. Moreover, we achieve second order in time, while our scheme also satisfies a discrete version of the balance law controlling the growth of the solution (see Propositions 2.1 and 2.2).
1.2. Time semi-discrete scheme
The scheme we introduce for the simulation of equation (1.1) can be described as a relaxation-Crank-Nicolson in time, and is inspired by [3, 6, 7]. We start from equation (1.1) and write it in system form using the auxiliary variable
on equipped with periodic boundary conditions,
Note that the first equation in (1.3) is obtained by a linear extrapolation from the points and Moreover, the second equation in (1.3) is, essentially, a Crank-Nicolson time discretization of the second equation in (1.2). As expected, this leads to a second-order-in-time scheme. The auxiliary variable lives on a staggered timegrid Its introduction removes the need for solving a nonlinear equation in each timestep, but requires the additional initialization of for the first timestep. This could be done either in the straightforward way or by solving an auxiliary problem – cf. Section 3.2 for more details.
2. The balance laws
If the background is zero, then equation (1.1) conserves the norm, . It is easy to check that the numerical scheme inherits this property automatically. However, for general the norm of the solution can grow – and this is the main phenomenon of interest. In such non-conservative problems, where the solution can grow hundreds of times in size, it is vital to have a guarantee that the rate of growth is qualitatively correct.
An norm balance law for the general continuous problem relates the rate of growth of the solution with the data of the problem. Our numerical scheme inherits this balance law at the discrete level.
On the continuous level, the balance law is as follows:
Proposition 2.1.
Consider equation (1.1), assume that is a smooth function, and that the given autocorrelation function is where is a nonnegative smooth function of compact support.
Then the potential is real valued for all times. Moreover, if we denote it follows that
Equation (2.1) holds when and exhibits rapid decay, or when and satisfies periodic boundary conditions.
Proof: First of all recall that the initial condition for (1.1) is Hermitian, This Hermitianity is preserved in time, and means that is real valued for all times. The real-valuedness of the potential is crucial in what follows.
For the well-posedness of equation (1.1) on full space, see [2]. In particular, for smooth initial data of rapid decay, the solution remains smooth and with rapid decay. For the adaptation to periodic boundary conditions, see [5]. Thus, all the integrals appearing below exist.
With regard to the growth of we readily observe that
The boundary conditions come into play only at the integration by parts of the Laplacian terms; the proof applies in the same way for full space with decay at infinity or if satisfies periodic boundary conditions. ∎
Essentially the same computation works out at the discrete level for the numerical scheme (1.3). We present it here in the time-discrete form for simplicity. The same carries over to the fully discrete level, where the discrete norm would appear in the place of the continuous norm.
Proposition 2.2.
Consider the time-discrete scheme (1.3), and denote
Then the following time-discrete balance law holds:
Proof: Multiplying equation (1.3) with and integrating in yields
By taking real parts the proof is complete. ∎
3. The numerical method
3.1. Fully discrete scheme
To implement the numerical scheme, finite differences are used on a uniform mesh on with spatial mesh size This makes easy the implementation of (which really is a trace operator). Moreover, it is important that the symmetry in is preserved in the spatial mesh. Let us denote by the array of the values of on the spatial mesh points and by the array of the values of on the spatial mesh.
Fourth-order central differences with periodic boundary conditions are used for the spatial discretization. The Kronecker product is used to create a sparse matrix implementing the hyperbolic Laplacian on a vectorized form of . Moreover, a diagonal matrix is created to implement the pointwise product of with The timestep of equation (1.3) is executed by first updating then solving for and finally extracting from it, as is standard for Crank-Nicolson-type schemes:
The matrices involved are sparse, and that should be heavily used in the solution of the linear systems. A spectral method could be used to implement the hyperbolic Laplacian, but this would make the matrix dense, and would lead to much slower implementation of the timestep, even if fewer points are needed.
3.2. Initialization
In order to execute the timestep of (3.1)-(3.3) for the first time, a value for is needed. In [6] a naive initialization approach is discussed, which adapted for this problem would amount to the following
As we will see, this naive initialization leads to the correct order of convergence for but loss of order for
In [3, 7, 18], ideas amounting to a order approximation of are discussed. In this paper we define
This advanced initialization leads to schemes of the expected order in both space and time, for both and To illustrate this in detail, we use an exact solution and examine the experimental order of convergence in space and time with each initialization. Details for the exact solution can be found in Section 3.3. In Tables 1 and 2 the expected order in space and time, respectively, is observed when the advanced initialization is used. In Table 3 it is seen that is clearly first order in time with the naive initialization. In Table 4 it is seen that with the naive initialization the spatial order for is degraded in practice, most likely because of the large time error. Note that denotes the spatial mesh size, while the timestep.
EOC for | EOC for | ||||
0.040019 | 0.021213 | 0.0035179 | 0.006808 | – | – |
0.040019 | 0.015 | 0.0017558 | 0.0034016 | 1.9696 | 1.9666 |
0.040019 | 0.010607 | 0.00087194 | 0.0016924 | 2.0005 | 1.9952 |
0.040019 | 0.0075 | 0.00043747 | 0.00084909 | 2.0093 | 2.0094 |
EOC for | EOC for | ||||
0.28243 | 0.0005 | 0.0048766 | 0.0081436 | – | – |
0.19917 | 0.0005 | 0.0012492 | 0.0020618 | 3.8994 | 3.9329 |
0.14064 | 0.0005 | 0.00031432 | 0.00052044 | 3.9655 | 3.9563 |
0.099584 | 0.0005 | 8.1076e-05 | 0.00013488 | 3.9256 | 3.9117 |
3.3. Experimental order of convergence
In the case of zero background, we can have exact solutions derived from exact solutions of the underlying nonlinear Schrödinger equation. Let and set in equation (1.1). If we choose a soliton amplitude and a soliton group velocity then setting and it follows that
is an exact solution of (1.1). Here we set For these choices of parameters, is numerically zero outside for Thus, when periodic boundary conditions are used on the exact solution can be periodized to
potentially allowing very long times to be considered. This is the exact solution on which Tables 1 through 4 are based, along with Figures 1 and 2. Observe that the solution has modulus of in value and derivatives, and this stays so for all times, making the interpretation of the errors straightforward.
The experimental order of convergence (EOC) values in row of each Table are derived from the runs in rows and The errors are defined as
with respect to the exact solution (3.7), where is the final time. Note that denotes the spatial mesh size, while the timestep.
EOC for | EOC for | ||||
0.040019 | 0.021213 | 0.0039656 | 0.063804 | – | – |
0.040019 | 0.015 | 0.0019807 | 0.045002 | 1.9675 | 0.98948 |
0.040019 | 0.010607 | 0.00098416 | 0.031765 | 1.999 | 0.99551 |
0.040019 | 0.0075 | 0.00049403 | 0.022433 | 2.0079 | 1.0134 |
EOC for | EOC for | ||||
0.28243 | 0.0005 | 0.0048769 | 0.0080754 | – | – |
0.19917 | 0.0005 | 0.0012493 | 0.0020243 | 3.8992 | 3.9613 |
0.14064 | 0.0005 | 0.00031454 | 0.0015111 | 3.9637 | 0.8402 |
0.099584 | 0.0005 | 8.1285e-05 | 0.0014965 | 3.9202 | 0.0282 |
The coarsest used is relatively large at but even so we get at least two significant digits correct in all cases, for both and
With regard to the naive initialization, Table 3 shows that is clearly first order in time. Extrapolating from Table 3, it follows that even for the time error for is expected to be Indeed, in the last two rows of Table 4, we see that the error in does not drop below when and Thus, the fourth order of convergence in space for which is initially observed for relatively large spatial mesh sizes (first two rows of Table 4), in practice is degraded by the large time error for smaller spatial mesh sizes.
3.4. Qualitative behavior of the error
Some observations can be made for the qualitative behavior of the error. These are based on a relatively long time run with the soliton-type solution of (3.7). The soliton exits from the top right corner of the computational domain, and re-enters from the bottom left corner; a fine discretization is used for this run, see Figure 1.
It is seen that the error in is not very sensitive to the error in the latter grows faster and to larger values. There is some kind of averaging taking place, and understanding that could be vital in developing optimal theory for this problem.
Another interesting phenomenon is that, after a rapid initial growth, the error seems to grow roughly linearly in time (and not exponentially, despite the problem being nonlinear).
Finally, in Figure 2 we see that the main part of the error in this problem is in fact due to the phase of the numerical solution staying slightly behind the exact solution, while the shape of the soliton is preserved to very high accuracy.
It should be noted that the continuous problem (1.1) is known to have several invariants, namely
i.e. for These first appeared in [14], and are derived in our notation for completeness in Lemma A.1 in the Appendix. They are based on the symmetries of and stay constant even if grows by factors of hundreds (i.e. they involve large cancellations in general). They can be used as a diagnostic to gauge the quality of the numerical solution. and can be preserved to 2 decimal digits in most computations. involves second order derivatives of the solution, and can be hard to preserve to good accuracy when the solution grows hundreds of times in size. Nevertheless, all invariants are preserved to at least two digits in all the runs presented in Section 3.
4. Landau damping and modulation instability
4.1. The bifurcation
A central feature of the Alber equation (1.1) is the bifurcation between Landau damping and modulation instability. In the stable case, the inhomogeneity disperses despite the presence of the homogeneous background which would be expected to act as a source term. On the other hand, in the unstable case, the inhomogeneity triggers unstable modes that grow exponentially – at least initially.
The linear instability condition is formulated in terms of the background power spectrum of the problem, [1]. The sufficient condition for instability is [1, 2]
where is the Hilbert transform and is the divided difference of the background spectrum with increment
We will discuss constructive ways to check (4.1) in the end of Section 4.1.
If (4.1) holds for some then that is an unstable wavenumber. The instability is a baseband modulation instability in the sense that unstable wavenumbers (when they exist) are of the form i.e., large enough wavenumbers are always stable. Moreover, the bandwidth of unstable wavenumbers, controls the lengthscales in the emerging coherent structure (larger bandwidth leads to more localized coherent structures), while the largest controls the rate of growth of the inhomogeneity, according to the linear stability analysis [4]. An important question is whether the fully nonlinear dynamics stay close to the linearized dynamics, especially for physically realistic inhomogeneities. For example, many nonlinear Landau damping results have extremely stringent smallness conditions for the inhomogeneity and its derivatives up to high order, which are questionable in any real-world system.
Another question that is hotly debated in the ocean waves community is the nature of the transition from stable to unstable behavior, and especially how it manifests phenomenologically in the ocean waves context. To fix ideas, consider the one-parameter family of spectra of the form for some smooth, unimodal profile It is easy to see that there exists a critical value so that if the spectrum is stable (in the sense of condition (4.1)), while for it is unstable. So the question is, how different are the solutions of (1.1) for versus Is there a striking change in the qualitative behavior of solutions around
A key piece of context for this question is the timescales of interest: physically, we expect the original equation (1.1) to be valid for an timescale but not for In that context, instability that takes an asymptotically long time to manifest, would appear similar to stability when [4]. In Section 4.2 we investigate numerically how well this asymptotic prediction of slow growth of the instability holds up in the fully nonlinear simulation.
Another aspect of this question is how large the inhomogeneity eventually becomes. In the ocean wave context, is compared to an homogeneous background. Thus, even if itself changes a lot, if it remains dominated by the background then the phenomenology of the sea state does not change noticeably. In this context the amplification factor is relevant: how much larger than the background can the inhomogeneity become? An amplification factor of at least is needed in order for the instability to produce dominant effects. This question is investigated in Section 4.2 for the first time by solving the fully nonlinear equation (1.1). In particular, this question cannot be studied through the linear stability analysis.
A subtle issue in this discussion is the localization of the problem from full space to a finite domain The linear stability analysis is typically carried out on full space. In [5] the question of how to localize the Alber equation and its linear stability analysis on a finite computational domain was investigated. It was found that setting
leads to the same stability condition as full space, when is large enough. In other words, we start from the continuous power spectrum (often measured in the field in the ocean waves context), and proceed to generate an autocorrelation function according to (4.3). This in turn provides a benchmark for the size of required: as long as the effective support of is well contained in then numerically and the of (4.3) are interchangeable.
The original instability condition (4.1) is not straightforward to check, as it involves the existence or not of solutions in a system of two equations (real and imaginary parts of ) in three unknowns (). A constructive way to check stability was introduced in [2]; the curve
is drawn on the complex plane. This is a closed curve starting and ending at (and it might be self-intersecting for multimodal spectra ). If the winding number of around is nonzero, then the wavenumber is unstable. This method computes the solutions of (4.1), but it only clarifies the existence of – not their values. In Figure 3 this technique is demonstrated for the Gaussian spectrum (4.5) with From this, the bandwidth of unstable wavenumbers can be computed. For the required computation of the Hilbert transform we use Weideman’s rational eigenfunction method [17].
Finally, it is worth noting that, in the ocean waves problem, the potential controls how large the standard deviation of the sea state from the homogeneous background is at location and time Thus, an appropriate way to keep track of the size of the inhomogeneity over space and time is through
4.2. Numerical investigation
The onset of modulation instability is not fully understood yet, but it is considered one of the possible mechanisms behind the growth of oceanic rogue waves [9]. Hence investigating its fully nonlinear behavior presents great interest both from a theoretical dynamical systems point of view, as well as an ocean waves point of view. Capturing the growth (or not) of the inhomogeneity is precisely why the discrete balance law of Proposition 2.2 is important, as it guarantees that the numerical solution qualitatively follows the exact one. Moreover, the second order in time and fourth order in space of the numerical method presented here makes tractable the large-domain and long-time computations required to conclusively investigate nonlinear Landau damping (in the stable case), and the growth of coherent structures (in the unstable case).
For clarity and compactness, here we investigate problems where all the terms are of comparable sizes, and use a Gaussian power spectrum of
The parameter controls the RMS wave amplitude in the background sea state. Moreover, this is a relatively narrow power spectrum, i.e., the situation in which modulation instability is expected to appear. The initial inhomogeneity is of the form
In order to have directly comparable results for different values of we used
in all runs. Values drawn from a complex normal distribution for yield similar results, i.e., the results are not particularly sensitive to the exact values. The strongest dependence of the results on the initial data is through the overall size of the initial condition, as for smaller initial inhomogeneities it can take substantially longer for the inhomogeneity to grow in the unstable case. The initial datum of (4.6), (4.7) has and in other words it is small but not vanishingly so. We use an initial inhomogeneity that is not symmetric nor a very simple function to highlight the robustness of the qualitative behavior of the problem. All numerical results in this Section, including Figures 4-7, are carried out with as in (4.6), (4.7), on a computational domain with timestep size and space mesh size
In the case of modulation instability, the inhomogeneity first grows exponentially to a localized coherent structure with values of and then a distinctive pattern emerges, of similar localized structures spreading outwards in space to form a lattice within a space-time cone. A typical example of the space-time lattice of coherent structures can be seen in Figure 4. This matches very closely the universal behavior of modulation instability in other wave problems [8].
When this is a relatively intense background with giving rise to a relatively strong instability. Here the values of on the peaks are clealry larger than the background, i.e. they describe dominant effects. The two distinct stages are visible in Figures 4 and 5: first, there is exponential growth of the inhomogeneity to a single, localized coherent structure. Once the first coherent structure is formed, there is no more growth in the norm, but the pattern propagates outwards forming the lattice-within-a-cone.
For purposes of comparison with the linear stability analysis of the problem, we also solve the linearized Alber equation,
In Figure 5 we see the evolution in time of the and norms for and for By or so the solutions of both the fully nonlinear and linearized equations have hit their maximum rate of growth. This picking-up-steam stage can take much longer when the instability is weaker (e.g., smaller ). Figure 5 shows that the linearized equation (4.8) captures very well the initial stage of exponential growth of the unstable modes. On the other hand, the linearized equation completely misses the formation of the space-time cone and the maxing out of the norm. In conclusion, the linear stability analysis is found to accurately predict the onset of modulation instability, as well the rate of growth in the early stage. However, it clearly misses things like the maximum amplification factor , the space-time cone etc. Full solution of the nonlinear equation (1.1) is required for that.
In Figure 6 we can see how the qualitative behavior changes with the strength of the background The values also lead to instability, but one that grows more slowly and to a smaller maximum, as compared to Perhaps more importantly than the raw amplitude of the inhomogeneity, the amplification factors are also noticeably smaller. As can be seen in Table 5, below the observed maximum amplitude of is smaller than the background In that sense, even if technically instability is present (and indeed the inhomogeneity grows by at least one order of magnitude from its initial size), it does not describe a dominant effect when compared to the homogeneous background.
One definition of oceanic rogue waves is waves that have heights at least twice as large as the RMS wave height of the sea state – i.e. an amplification factor of at least in practice probably closer to In that sense, only strong enough instabilities can lead to rogue waves. It is remarkable that within a small range the amplification factor can jump from to This is a sudden change in the qualitative behavior of the sea state. However, it does not happen at the onset of linear instability, but instead when the instability becomes strong enough.
5. Discussion and Outlook
A relaxation-Crank-Nicolson scheme for the von Neumann equation with power nonlinearity was introduced. The scheme is second-order in time, and requires the solution of a linear system in each timestep, i.e., it is explicit in the nonlinearity. This is achieved by the introduction of an auxiliary variable on a staggered time-grid. The scheme is implemented with fourth-order finite differences in space. It inherits at the discrete level the problem’s balance law (Proposition 2.2), making it particularly well suited to investigate the possible growth of the solution.
In Section 3 the scheme is validated with a soliton-derived exact solution. It is found that the initialization of the auxiliary variable is important so that the scheme achieves the expected order in both space and time, for both the main and auxiliary variables. Moreover, it is found that the main variable is to a surprising extent not affected by the error in the auxiliary variable, for example exhibiting the correct order of convergence even if the auxiliary variable does not (Tables 3 and 4). This seems to indicate some kind of averaging, and understanding it fully could be vital in fully analyzing this kind of scheme. Finally, in Figure 2 the bulk of the error is seen to be due to error in phase velocity, at least for this soliton-like solution.
In Section 4 the bifurcation between Landau damping and modulation instability is discussed and numerically investigated. The universal space-time cone [8] is recovered when modulation instability is present. The prediction that, in borderline unstable cases, the inhomogeneity grows very slowly [4] is confirmed. Moreover, when the instability is weak, the inhomogeneity maxes out to small amplitudes. Perhaps more importantly for ocean waves, weak instabilities lead to small amplification factors, i.e. the inhomogeneity remains dominated by the homogeneous background in the fully nonlinear dynamics. On the other hand, strong enough instabilities do lead to dominant effects compared to the homogeneous background. Finally, if the homogeneous background becomes smaller than a threshold ( in this case), there is no modulation instability at all. Instead, nonlinear Landau damping is observed – a dispersion-dominated behaviour, where the linearized Alber equation (4.8) stays very close to the fully nonlinear dynamics of (1.1).
In the context of ocean waves where this equation originates [1], the scaling of the problem would be Modulation instability is indeed expected in the most intense sea states in the North Atlantic, [2]. However, because of the aforementioned scaling of the problem, long times and large domains would be required to simulate it where it is present. Using the numerical method introduced here, it is now possible for the first time to carry out a systematic study of the amplification factors, and their sensitivity to the size of the computational domain, initial condition etc. This is a fully nonlinear question, only accessible by simulation.
Appendix A Derivation of invariants
The following invariants for (1.1) were introduced in [14]. We include them here, along with their derivation in our notation, for completeness.
Lemma A.1.
Remark: The function is real-valued by the symmetry of the problem, but not non-negative. Thus the conservation of is compatible with potentially large values of but there will have to be positive and negative values so that the total integral is preserved.
Proof: To prove this lemma, it is helpful to use the change of variables
Originally are two position variables, and has the physical meaning of “mean position” while of “lag” between the two positions. For any observe that
In this formalism, equation (A.4) is elaborated to
Proof of equation (A.8): setting in equation (A.6) and integrating in leads to
by performing integration by parts in
Proof of equation (A.9): differentiating equation (A.7) with respect to yields
Now setting this becomes
Integrating over gives
Equation (A.9) follows by integration by parts in the two last terms.
Proof of equation (A.10): by differentiating twice in equation (A.7) becomes
By setting we get
by integrating in and some obvious integrations by parts
Moreover, by multiplying equation (A.7) with setting and integrating in we get
By combining equations (A.12) and (A.13), relation (A.10) follows. ∎