### v2

# A second-order-in-time scheme for the von Neumann equation with power nonlinearity and simulation of the onset of instability

Abstract.
The von Neumann equation with power nonlinearity serves as a statistical model for nonlinear waves, and it exhibits a bifurcation between stable and unstable regimes. In ocean waves it is known as the Alber equation, and its bifurcation is considered important for understanding rogue waves, a key problem in marine safety. Despite this, only one first-order-in-time numerical method exists in the literature for its simulation. In this paper, we introduce a structure-preserving, linearly implicit, second-order-in-time scheme for its numerical solution. We employ fourth-order finite differences in space, allowing for much larger and faster simulations. As an illustrative example, we explore the onset of modulation instability. We verify that the linear stability analysis accurately predicts the initial growth phase, but find that it fails to forecast the maximum amplitude or the formation of a coherent structure in the nonlinear regime. Monte Carlo simulations with Gaussian background spectra reveal that the maximum amplitude depends solely on the homogeneous background, and not on the initial inhomogeneity. For weak instabilities, the inhomogeneity grows substantially from its initial condition, but it remains small compared to the background. On the other hand, past a critical threshold, the instability grows to yield dominant effects. Thus, beyond the onset of linear instability, the solution of the fully nonlinear problem presents a second bifurcation – at the point which inhomogeneous features become dominant compared to the background.
Key words and phrases:
relaxation scheme, structure preserving, finite differences, Alber equation, modulation instability, nonlinear Landau damping.
1. Introduction
The von Neumann equation with power nonlinearity and homogeneous background, also known as the Alber equation, is given by
$\begin{array}[]{c}i\partial_{t}u+p(\Delta_{x}-\Delta_{y})u+q\Bigl(V(x,t)-V(y,t)\Bigr)\,\Bigl(\Gamma(x-y)+u(x,y,t)\Bigr)=0,\\[8.0pt] V(x,t)=u(x,x,t),\qquad u(x,y,0)=u_{0}(x,y),\end{array}$ (1.1)
where $p$ and $q$ are real-valued parameters, $\Gamma$ is a given autocorrelation function, and $u_{0}$ is Hermitian, $u_{0}(x,y)=\overline{u_{0}(y,x)}.$ The autocorrelation $\Gamma$ describes the known homogeneous background of the problem, and the unknown function $u$ represents the inhomogeneity.
Equation (1.1) appears in the context of ocean waves [1, 5, 13, 14, 21, 22, 23, 25, 26], and is known to exhibit a bifurcation between a stable regime (akin to nonlinear Landau damping) and an unstable regime (generalized modulation instability), depending on the homogeneous background $\Gamma.$ Recent theoretical works on its well-posedness include [11, 17, 24]. Related studies in other branches of physics include [12, 16, 18, 19].
In this paper, we introduce a second-order-in-time relaxation scheme, which inherits the problem’s key balance law at the discrete level. An advanced initialization for the scheme is introduced and validated against exact solutions. This is required to ensure that both the main and relaxation variables achieve the expected order of convergence. Finally, the scheme is used to investigate both modulation instability and nonlinear Landau damping for equation (1.1).
We are assuming the existence of a smooth solution. This is based on the results of [3, 6].
Sections 2 and 3 describe in a self-contained way the numerical method, its features and validation. This should be accessible to a broad audience, and to that end we strive to make minimal reference to the analytical and phenomenological properties of equation (1.1). On the other hand, in Section 4 we use the numerical method to investigate key questions about equation (1.1), such as the onset of instability, the maxima of the fully developed instability, and the dispersion of inhomogeneities in the stable case. Thus, Section 4 necessarily relies heavily on the state of the art for equation (1.1). The key finding of a second bifurcation is described in Section 4.3 and summarized in Figure 10.
1.1. Numerical methods for the von Neumann equation with power nonlinearity (Alber equation): state of the art
To the best of our knowledge, there are only two papers dealing systematically with the numerical solution of the Alber equation. These are [23, 26], and both treat equation (1.1) after a change of variables, to “mean position” $X=\frac{x+y}{2},$ and “lag” r=xyr=x-y (this change of variables first appears in [1], and it is also employed in the calculations of Appendix A). Both papers [23, 26] use a forward Euler solver for time-stepping, and second-order finite differences in space. Moreover, a non-standard boundary condition in the variable $r$ is used.
The works [23, 26] are fundamental for the field, and they highlight the challenges this non-standard equation poses. The introduction of boundary conditions in $r$ introduced in [26] attempt to match the far-field behaviour of the equation with a known background behaviour derived from a box spectrum. This is modified in [23] to match a JONSWAP spectrum. The idea behind these boundary conditions is that the inhomogeneity should be zero at  infinity, and it is a natural attempt to enforce the “zero at large rr” boundary condition for the inhomogeneity. As we see in the numerical results, in the case of modulationally unstable problems, the inhomogeneity extends its support, and inevitably reaches the boundary of the computational domain.
Our approach is to solve directly (1.1) with periodic boundary conditions in the $x,y$ variables. This preserves by construction the original physical context [20], and is the most widely used boundary condition with phase-resolved simulations for the same problem cf., e.g., [15, 21]. This also treats in a unified way any background spectrum (i.e. there is no need to specially fine tune the boundary conditions to each spectrum as in [23]). In any case, the high order of convergence of our method allows us to use sufficiently large computational domains so that the modulation instability is fully developed before any interaction with the boundary (cf., e.g., Figure 6).
Crucially, our scheme satisfies a discrete version of the balance law controlling the growth of the solution (see Propositions 2.1 and 2.2). This is particularly important for the unstable cases, when the solution $u$ grows by a factor of tens or hundreds.
1.2. Time semi-discrete scheme
The scheme we introduce for the simulation of equation (1.1) can be described as a relaxation-Crank-Nicolson in time, and is inspired by [4, 8, 9]. We start from equation (1.1) and write it in system form using the auxiliary variable $\phi$
[…]
on $\frac{L}{2},\frac{L}{2}]$ equipped with periodic boundary conditions,
$(x,y)\in[-\frac{L}{2},\frac{L}{2}]^{2},\quad\partial_{y}^{j}u(x,-\frac{L}{2})=\partial_{y}^{j}u(x,\frac{L}{2}),\,\,\partial_{x}^{j}u(-\frac{L}{2},y)=\partial_{x}^{j}u(\frac{L}{2},y),\quad j\in\{0,1\}.$
Now equation (1.2) is discretized in time to
[…]
where
$U^{n+\frac{1}{2}}=U^{n+\frac{1}{2}}(x,y):=\frac{U^{n+1}(x,y)+U^{n}(x,y)}{2}.$ (1.4)
Note that the first equation in (1.3) is obtained by a linear extrapolation from the points $t^{n}$ and $t^{n+\frac{1}{2}}.$ Moreover, the second equation in (1.3) is, essentially, a Crank-Nicolson time discretization of the second equation in (1.2). As expected, this leads to a second-order-in-time scheme. The auxiliary variable $\Phi^{n+\frac{1}{2}}$ lives on a staggered timegrid $t^{n+\frac{1}{2}}:=(t^{n+1}+t^{n})/2.$ Its introduction removes the need for solving a nonlinear equation in each timestep, while keeping strong stability features – analogous to those of the fully implicit Crank-Nicolson for the nonlinear Schrödinger equation. The price for that is requiring the additional initialization of $\Phi^{-\frac{1}{2}}$ for the first timestep. This question requires some nuance, and it is discussed in detail in Section 3.2 for more details.
2. The balance laws
If the background $\Gamma$ is zero, then equation (1.1) conserves the $L^{2}$ norm, $\|u(t)\|_{L^{2}_{x,y}}=\|u(0)\|_{L^{2}_{x,y}}$. It is easy to check that the numerical scheme inherits this property automatically. However, for general $\Gamma,$ the $L^{2}$ norm of the solution can grow – and this is the main phenomenon of interest. In such non-conservative problems, where the solution can grow hundreds of times in size, it is vital to have a guarantee that the rate of growth is qualitatively correct.
An $L^{2}$ norm balance law for the general continuous problem relates the rate of growth of the solution with the data of the problem. Our numerical scheme inherits this balance law at the discrete level.
On the continuous level, the balance law is as follows:
Proposition 2.1.
Consider equation (1.1), assume that $u_{0}$ is a smooth function, and that the given autocorrelation function $\Gamma(x)$ is $\Gamma(x)=\mathcal{F}^{-1}_{k\to x}[P(k)],$ where $P(k)$ is a nonnegative smooth function of compact support.
Then the potential $V(x,t)=u(x,x,t)$ is real valued for all times. Moreover, if we denote $\mathcal{M}(t):=\|u(t)\|_{L^{2}}^{2},$ it follows that
$\frac{d}{dt}\mathcal{M}(t)=2\mathrm{Re}\Bigl[iq\int\limits_{x,y}\Bigl(V(x,x,t)-V(y,y,t)\Bigr)\Gamma(x-y)\overline{u(x,y,t)}dxdy\Bigr].$ (2.1)
Equation (2.1) holds when $(x,y)\in\mathbb{R}^{2n}$ and $u_{0}$ exhibits rapid decay, or when $(x,y)\in[-\frac{L}{2},\frac{L}{2}]^{2n}$ and $u$ satisfies periodic boundary conditions.
Proof: First of all recall that the initial condition $u_{0}(x,y)$ for (1.1) is Hermitian, $u_{0}(x,y)=\overline{u_{0}(y,x)}.$ This Hermitianity is preserved in time, and means that $V(x,t)=u(x,x,t)$ is real valued for all times. The real-valuedness of the potential is crucial in what follows.
For the well-posedness of equation (1.1) on full space, $(x,y)\in\mathbb{R}^{2n},$ see [3]. In particular, for smooth initial data of rapid decay, the solution remains smooth and with rapid decay. For the adaptation to periodic boundary conditions, see [6]. Thus, all the integrals appearing below exist.
With regard to the growth of $\mathcal{M}(t),$ we readily observe that
$\begin{array}[]{l}\frac{d}{dt}\mathcal{M}(t)=\frac{d}{dt}\langle u,u\rangle=2\mathrm{Re}\Bigl[\langle u_{t},u\rangle\Bigr]=\\ =2\mathrm{Re}\Bigl[\Bigl\langle ip(\Delta_{x}-\Delta_{y})u+iq\Bigl(V(x,x,t)-V(y,y,t)\Bigr)\,\Bigl(\Gamma(x-y)+u\Bigr)\,,\,\,u\Bigr\rangle\Bigr]=\\ =2\mathrm{Re}\Bigl[-ip\Bigl\langle\nabla_{x}u,\nabla_{x}u\Bigr\rangle+ip\Bigl\langle\nabla_{y}u,\nabla_{y}u\Bigr\rangle+\hfill\\ \hfill+\Bigl\langle iq\Bigl(V(x,x,t)-V(y,y,t)\Bigr)\,\Bigl(\Gamma(x-y)+u\Bigr)\,,\,\,u\Bigr\rangle\Bigr]=\\ =2\mathrm{Re}\Bigl[iq\int\limits_{x,y}\Bigl(V(x,x,t)-V(y,y,t)\Bigr)\Gamma(x-y)\overline{u(x,y,t)}dxdy\Bigr].\end{array}$
The boundary conditions come into play only at the integration by parts of the Laplacian terms; the proof applies in the same way for full space $\mathbb{R}^{2n}$ with decay at infinity or if $u$ satisfies periodic boundary conditions. ∎
Essentially the same computation works out at the discrete level for the numerical scheme (1.3). We present it here in the time-discrete form for simplicity. The same carries over to the fully discrete level, where the discrete $L^{2}$ norm would appear in the place of the continuous $L^{2}$ norm.
Proposition 2.2.
Consider the time-discrete scheme (1.3), and denote
$\mathcal{M}^{n}:=\|U^{n}\|_{L^{2}}^{2}.$ (2.2)
Then the following time-discrete balance law holds:
$\frac{\mathcal{M}^{n+1}-\mathcal{M}^{n}}{\tau}=2\mathrm{Re}\Bigl[iq\Bigl\langle\Phi^{n+\frac{1}{2}}\Gamma(x-y),U^{n+\frac{1}{2}}\Bigr\rangle\Bigr].$ (2.3)
Remark: Equation (2.3) is a time-discrete version of equation (2.1).
Proof: Multiplying equation (1.3) with $\overline{U^{n+\frac{1}{2}}}$ and integrating in $x,y$ yields
$\begin{array}[]{c}\langle U^{n+1}-U^{n},U^{n+\frac{1}{2}}\rangle=\langle i\tau p(\Delta_{x}-\Delta_{y})U^{n+\frac{1}{2}}+i\tau q\Phi^{n+\frac{1}{2}}(\Gamma(x-y)+U^{n+\frac{1}{2}}),U^{n+\frac{1}{2}}\rangle=\\[7.0pt] =i\tau p\langle\Delta_{x}U^{n+\frac{1}{2}},U^{n+\frac{1}{2}}\rangle-i\tau p\langle\Delta_{y}U^{n+\frac{1}{2}},U^{n+\frac{1}{2}}\rangle+\hfill\\ \hfill+i\tau q\langle\Phi^{n+\frac{1}{2}}\Gamma(x-y),U^{n+\frac{1}{2}}\rangle+i\tau q\langle\Phi^{n+\frac{1}{2}}U^{n+\frac{1}{2}},U^{n+\frac{1}{2}}\rangle.\end{array}$
By taking real parts the proof is complete. ∎
3. The numerical method
3.1. Fully discrete scheme
To implement the numerical scheme, finite differences are used on a uniform $N\times N$ mesh on $\frac{L}{2},\frac{L}{2}]^{2}$ with spatial mesh size $h,$ $Nh=L.$ This makes easy the implementation of $U(x,x,t)$ (which really is a trace operator). Moreover, it is important that the symmetry in $x,y$ is preserved in the spatial mesh. Let us denote by ${U}_{i,j}^{n}$ the array of the values of $U^{n}$ on the spatial mesh points $x_{i},y_{j},$ and by ${\Phi}_{i,j}^{n+\frac{1}{2}}$ the array of the values of $\Phi^{n+\frac{1}{2}}$ on the spatial mesh.
Fourth-order central differences with periodic boundary conditions are used for the spatial discretization. The Kronecker product is used to create a sparse $N^{2}\times N^{2}$ matrix $\mathfrak{D}_{H}$ implementing the hyperbolic Laplacian $\Delta_{x}-\Delta_{y}$ on a vectorized form of $U^{n+\frac{1}{2}}_{i,j}$. Moreover, a $N^{2}\times N^{2}$ diagonal matrix $\mathfrak{F}^{n+\frac{1}{2}}=\texttt{diag}(\texttt{vec}({\Phi}_{i,j}^{n+\frac{1}{2}}))$ is created to implement the pointwise product of $\Phi^{n+\frac{1}{2}}$ with $U^{n+\frac{1}{2}}.$ The timestep of equation (1.3) is executed by first updating $\Phi^{n+\frac{1}{2}}_{i,j},$ then solving for $U^{n+\frac{1}{2}}_{i,j},$ and finally extracting $U^{n+{1}}_{i,j}$ from it, as is standard for Crank-Nicolson-type schemes:
$\Phi^{n+\frac{1}{2}}_{i,j}=2\Bigl(U^{n}_{i,i}-U^{n}_{j,j}\Bigr)-\Phi^{n-\frac{1}{2}}_{i,j},$ (3.1)
$\Bigl(I-i\frac{p\tau}{2}\mathfrak{D}_{H}-i\frac{q\tau}{2}\mathfrak{F}^{n+\frac{1}{2}}\Bigr)\texttt{vec}(U^{n+\frac{1}{2}}_{i,j})=\texttt{vec}\bigl(U^{n}_{i,j}+i\frac{q\tau}{2}\Gamma(x_{i}-y_{j})\Phi^{n+\frac{1}{2}}_{i,j}\bigr),$ (3.2)
$U^{n+1}_{i,j}=2U^{n+\frac{1}{2}}_{i,j}-U^{n}_{i,j}.$ (3.3)
The matrices involved are sparse, and that should be heavily used in the solution of the linear systems. A spectral method could be used to implement the hyperbolic Laplacian, but this would make the matrix dense, and in many cases would lead to much slower implementation of each timestep, even if fewer points are needed.
3.2. Initialization
In order to execute the timestep of (3.1)-(3.3) for the first time, a value for $\Phi^{-\frac{1}{2}}_{i,j}$ is needed. In [8] a naive initialization approach is discussed, which adapted for this problem would amount to the following
$\mbox{Naive initialization: }\quad\Phi^{-\frac{1}{2}}_{i,j}=U^{0}_{i,i}-U^{0}_{j,j}.$ (3.4)
As we will see, this naive initialization yields the expected second order of convergence for $u,$ but results in first order convergence in time for $\phi.$
In [4, 9, 29], ideas amounting to a $2^{nd}$ order approximation of $\Phi(-\tau/2)$ are discussed. In this paper we define
$\begin{array}[]{c}\left\{\begin{array}[]{l}\Phi^{*}_{i,j}=U^{0}_{i,i}-U^{0}_{j,j},\\ \Bigl(I+i\frac{p\tau}{4}\mathfrak{D}_{H}+i\frac{q\tau}{4}\mathfrak{F}^{*}\Bigr)\texttt{vec}(U^{-\frac{1}{4}}_{i,j})=\texttt{vec}\bigl(U^{0}_{i,j}-i\frac{q\tau}{4}\Gamma(x_{i}-y_{j})\Phi^{*}_{i,j}\bigr),\\ U^{-\frac{1}{2}}_{i,j}=2U^{-\frac{1}{4}}_{i,j}-U^{0}_{i,j},\end{array}\right\}\\[18.0pt] \mbox{Advanced initialization: }\quad\Phi^{-\frac{1}{2}}_{i,j}=U^{-\frac{1}{2}}_{i,i}-U^{-\frac{1}{2}}_{j,j}.\par\end{array}$ (3.5)
This advanced initialization leads to schemes of the expected order in both space and time, for both $u$ and $\phi.$ To illustrate this in detail, we use an exact solution and examine the experimental order of convergence in space and time with each initialization. Details for the exact solution can be found in Section 3.3.
Table 1. Time EOC for $u,\phi;$ advanced initialization. We denote the spatial mesh size by $h$ and the timestep by $\tau.$
$h$ $\tau$ $\mathcal{E}_{u}$ $\mathcal{E}_{\phi}$ EOC for $u$ EOC for $\phi$
0.04 0.03 0.007135 0.01397
0.04 0.02121 0.003517 0.006808 2.078 2.113
0.04 0.015 0.001755 0.003401 1.969 1.966
0.04 0.0106 0.0008719 0.001692 2.000 1.995
Table 2. Spatial EOC for $u,\phi;$ advanced initialization. We denote the spatial mesh size by $h$ and the timestep by $\tau.$
$h$ $\tau$ $\mathcal{E}_{u}$ $\mathcal{E}_{\phi}$ EOC for $u$ EOC for $\phi$
0.4006 0.0005 0.01997 0.03305
0.2824 0.0005 0.004876 0.008143 4.032 4.006
0.1991 0.0005 0.001249 0.002061 3.899 3.932
0.1406 0.0005 0.0003143 0.0005204 3.965 3.956
In Tables 1 and 2 the expected order in space and time, respectively, is observed when the advanced initialization is used. In Table 3 it is seen that $\phi$ is clearly first order in time with the naive initialization. In Table 4 it is seen that with the naive initialization the spatial order for $\phi$ is degraded in practice, most likely because of the large time error. Note that $h$ denotes the spatial mesh size, while $\tau$ the timestep.
3.3. Experimental order of convergence
In the case of zero background, $\Gamma=0,$ we can have exact solutions derived from exact solutions of the underlying nonlinear Schrödinger equation. Let $\Gamma=0$ and set $p=1.7,$ $q=1.1$ in equation (1.1). If we choose a soliton amplitude $A\in\mathbb{R}$ and a soliton group velocity $v\in\mathbb{R},$ then setting k=v/(2p)k={v}/{(2p)} and $B=A\sqrt{{q}/{(2p)}},$ it follows that
$\begin{array}[]{c}u(x,y,t)=A^{2}\mathrm{sech}(B(x-vt))\mathrm{sech}(B(y-vt))\,e^{ik(x-y)}\end{array}$ (3.6)
is an exact solution of (1.1). Here we set $A=1.3,$ $v=3.1.$ For these choices of parameters, $u(x,y,0)$ is numerically zero outside $\Omega:=[-\frac{L}{2},\frac{L}{2}]^{2}$ for $L=10\pi/k\approx 34.45.$ Thus, when periodic boundary conditions are used on $\Omega,$ the exact solution can be periodized to
$\begin{array}[]{c}w(\xi)=(\xi+\frac{L}{2}\mod L)\,\,-\frac{L}{2}\\ u(x,y,t)=A^{2}\mathrm{sech}\bigl(Bw(x-vt)\bigr)\mathrm{sech}\bigl(Bw(y-vt)\bigr)\,e^{ik(x-y)}\end{array}$ (3.7)
potentially allowing very long times to be considered. This is the exact solution on which Tables 1 through 4 are based, along with Figures 1 and 2. Observe that the solution has modulus of $O(1)$ in value and derivatives, and this stays so for all times, making the interpretation of the errors straightforward.
The experimental order of convergence (EOC) values in row $i$ of each Table are derived from the runs in rows $i$ and i1.i-1. The errors are defined as
$\mathcal{E}_{u}:=\max\limits_{i,j}\left|U^{N}_{i,j}-u(x_{i},y_{j},t^{N})\right|$ (3.8)
$\mathcal{E}_{\phi}:=\max\limits_{i,j}\left|\Phi^{N-\frac{1}{2}}_{i,j}-\phi(x_{i},y_{j},t^{N-\frac{1}{2}})\right|$ (3.9)
with respect to the exact solution (3.7), where $t^{N}=T\approx 1.1$ is the final time. Note that $h$ denotes the spatial mesh size, while $\tau$ the timestep.
Table 3. Time EOC for $u,\phi;$ naive initialization. We denote the spatial mesh size by $h$ and the timestep by $\tau.$
$h$ $\tau$ $\mathcal{E}_{u}$ $\mathcal{E}_{\phi}$ EOC for $u$ EOC for $\phi$
0.04 0.03 0.008034 0.08834
0.04 0.02121 0.003965 0.0638 2.074 0.9561
0.04 0.015 0.001980 0.045 1.967 0.9894
0.04 0.0106 0.0009841 0.03176 1.999 0.9955
Table 4. Spatial EOC for $u,\phi;$ naive initialization. We denote the spatial mesh size by $h$ and the timestep by $\tau.$
$h$ $\tau$ $\mathcal{E}_{u}$ $\mathcal{E}_{\phi}$ EOC for $u$ EOC for $\phi$
0.4006 0.0005 0.01997 0.03299
0.2824 0.0005 0.004876 0.008075 4.032 4.025
0.1991 0.0005 0.001249 0.002024 3.899 3.961
0.1406 0.0005 0.0003145 0.001511 3.963 0.8402
The coarsest $h$ used is relatively large at $h\approx 0.4,$ but even so we get two significant digits correct in both $u$ and $\phi$ in all cases.
With regard to the naive initialization, Table 3 shows that $\phi$ is clearly first order in time. Extrapolating from Table 3, it follows that even for $\tau=0.005,$ the time error for $\phi$ is expected to be $\mathcal{E}_{\phi}\approx 0.00146.$ Indeed, in the last two rows of Table 4, we see that the error in $\phi$ does not drop below $\mathcal{E}_{\phi}\approx 0.0015$ when $h=0.1406$ and $\tau=0.0005.$ Thus, the fourth order of convergence in space for $\phi,$ which is initially observed for relatively large spatial mesh sizes (first three rows of Table 4), in practice is degraded by the large time error for smaller spatial mesh sizes.
Figure 1. Evolution in time for the $l^{\infty}_{i,j}$ error for $u$ and $\phi,$ for the exact solution (3.7). The final time $T=11.1149$ used here amounts to a full lap of the computational domain of length $L=34.4562$, i.e. the soliton returns near the starting position. For this computation the mesh sizes used are $\tau=0.002$ and $h=0.12.$
3.4. Qualitative behavior of the error
Figure 2. Zoom in of the error in Figure 1 at the initial stage. The oscillations relax after some time. The naive initialization produces substantially more pronounced oscillations.
Figure 3. Constraint error, as defined in equation (3.10), for the example of Figure 1.
Some observations can be made for the qualitative behavior of the error. These are based on a relatively long time run with the soliton-type solution of (3.7), with the advanced initialization and well-resolved discretization, cf. Figure 1. The soliton exits from the top right corner of the computational domain, and re-enters from the bottom left corner, until it returns to the origin. The timestep is $\tau=0.002$ and the spatial mesh size $h=0.12.$
It is seen that the error in $U$ is not very sensitive to the error in $\Phi:$ the latter grows faster and to larger values. There is some kind of averaging taking place, and understanding that could be vital in developing optimal theory for this problem.
Another interesting phenomenon is that, after a rapid initial growth, the error seems to grow roughly linearly in time (and not exponentially, despite the problem being nonlinear).
If we look at more fine behaviours of the error, we see the trace of the initialization used in the oscillations of the exact error, cf. Figure 2. To the extent that smaller oscillations can be seen as better behavior of the scheme, this aspect also favors the advanced initialization over the naive one.
This scheme rests crucially on the introduction of an auxiliary variable on a staggered time-grid, $\Phi^{n+\frac{1}{2}}$ which aims to represent $u(x,x,t^{n+\frac{1}{2}})-u(y,y,t^{n+\frac{1}{2}}).$ The direct relationship between $\Phi^{n+\frac{1}{2}}$ and $U^{n}$ is through the first equation in (1.3), i.e. the $U^{n}$ interpolates appropriately in time the $\Phi^{n\pm\frac{1}{2}}.$ To see how well the relationship between the main and auxiliary variables holds up, one could ask how well the reverse holds, i.e. how well $\Phi^{n+\frac{1}{2}}$ interpolates $U^{n},$ $U^{n+1}.$ This is measured by the quantity
$\mathcal{E}_{constr}^{n}:=\max\limits_{i,j}\left|\frac{1}{2}\left(U^{n+1}_{i,i}-U^{n+1}_{j,j}-U^{n}_{i,i}+U^{n}_{j,j}\right)-\Phi^{n+\frac{1}{2}}_{i,j}\right|.$ (3.10)
To the best of our knowledge such measures of consistency for auxiliary variables do not have a standard name, and we call this the “constraint error”. This seems to fluctuate initially (roughly for the same time period as the oscillations of Figure 2), and then settles to a constant value, much smaller than the error – see Figure 3. Near the final time, when the soliton returns near its original position, we see a “refocusing” of the numerical artifacts – most clearly in Figure 3, but also some oscillations in the error in $\Phi$ in Figure 1. This is a consequence of periodic boundary conditions.
The apparent boundedness of this constraint error in time is a promising feature for the use of the scheme in long times, and is in contrast with other auxiliary variable kinds of schemes (such as the IEQ method for parabolic problems) where analogous quantities are known to grow [2].
Finally, by looking at the full difference between the numerical and exact solutions, we can get a feeling on where the bulk of the error comes from. In Figure 2 we see that the main part of the error in this problem is in fact due to the phase of the numerical solution staying slightly behind the exact solution, while the shape of the soliton is preserved to very high accuracy.
Figure 4. This plot is the real part of $u(t^{N})-U^{N}$ over $(x,y)\in[-\frac{L}{2},\frac{L}{2}]^{2}$ at final time, for the run described in Figure 1. The shape of the soliton is very well preserved; the bulk of the error is due to the phase of the numerical solution staying slightly behind that of the exact solution.
It should be noted that the continuous problem (1.1) is known to have several invariants, namely
$I_{1}[u(t)]:=\int\limits_{x}u(x,y,t)\Big|_{y=x}dx,$ (3.11)
$I_{2}[u(t)]:=\int\limits_{x}(\partial_{x}-\partial_{y})u(x,y,t)\Big|_{y=x}dx,$ (3.12)
$I_{3}[u(t)]:=\left(\frac{q}{p}\int\limits_{x}u^{2}(x,y,t)\Big|_{y=x}dx+\int\limits_{x}(\partial_{x}-\partial_{y})^{2}u(x,y,t)\Big|_{y=x}dx\right),$ (3.13)
i.e. $\frac{d}{dt}I_{j}[u(t)]=0$ for $j\in\{1,2,3\}.$ These first appeared in [26], and are derived in our notation for completeness in Lemma A.1 in the Appendix. They are based on the symmetries of $u,$ and stay constant even if $u$ grows by factors of hundreds, as in the modulationally unstable examples we will see later (i.e. they involve large cancellations in general). They can be used as a diagnostic to gauge the quality of the numerical solution. $I_{1}$ and $I_{2}$ can be preserved to 2 decimal digits in most computations. $I_{3}$ involves second order derivatives of the solution, and can be hard to preserve to good accuracy when the solution grows hundreds of times in size. All invariants $I_{j},$ $j=1,2,3,$ are preserved very well in the example of Figure 1; see Table 5.
Table 5. Invariants $I_{j},$ $j\in\{1,2,3\}$ for the example of Figure 1. The relative errors are computed as $\left|(I_{j}[u_{0}]-I_{j}[u(T)])/I_{j}[u_{0}]\right|.$
$I_{1}[u_{0}]$ $I_{2}[u_{0}]$ $I_{3}[u_{0}]$ Rel. error $I_{1}$ Rel. error $I_{2}$ Rel. error $I_{3}$
$4.568$ $8.331i$ 15.19-15.19 $2.309\cdot 10^{-8}$ $2.781\cdot 10^{-8}$ $1.689\cdot 10^{-5}$
4. Landau damping and modulation instability
4.1. The bifurcation
A central feature of the Alber equation (1.1) is the bifurcation between a linearly stable and linearly unstable regime. The stable regime is similar to Landau damping [3], and the inhomogeneity $u$ disperses despite the presence of the homogeneous background $\Gamma$ which would be expected to act as a source term. In the unstable case, the inhomogeneity $u$ triggers unstable modes that grow exponentially – at least initially. This is often called modulation instability, in analogy with the standard modulation instability of the nonlinear Schrödinger equation (NLS) [7, 28].
The linear instability condition is formulated in terms of the background power spectrum of the problem, $P(k)=\mathcal{F}_{x\to k}[\Gamma(x)]$ [1], and it is a Penrose-type condition. The sufficient condition for instability is [1, 3]
$\exists X_{*}\qquad\exists\omega_{*}\mbox{ with }\mathrm{Re}(\omega_{*})>0\qquad\mathbb{H}[D_{X_{*}}P](\omega_{*})=4\pi\frac{p}{q},$ (4.1)
where $\mathbb{H}$ is the Hilbert transform and $D_{X}P(k)$ is the divided difference of the background spectrum $P$ with increment $X,$
$D_{X}P(k)=\left\{\begin{array}[]{cc}\frac{P\left(k+\frac{X}{2}\right)-P\left(k-\frac{X}{2}\right)}{X},&X\neq 0,\\[6.0pt] P^{\prime}(k),&X=0.\end{array}\right$. (4.2)
We will discuss constructive ways to check condition (4.1) in the end of Section 4.1.
If (4.1) holds for some $X_{*},$ then that $X_{*}$ is an unstable wavenumber. The instability is a baseband modulation instability in the sense that unstable wavenumbers (when they exist) are of the form $|X_{*}|<X_{\text{max}},$ i.e., large enough wavenumbers are always stable. Moreover, the bandwidth of unstable wavenumbers, $2X_{\text{max}},$ controls the lengthscales in the emerging coherent structure (larger bandwidth leads to more localized coherent structures), while the largest $\mathrm{Re}(\omega_{*})$ controls the rate of growth of the inhomogeneity, according to the linear stability analysis [5]. An important question is whether the fully nonlinear dynamics stay close to the linearized dynamics, especially for physically realistic inhomogeneities. For example, many nonlinear Landau damping results have extremely stringent smallness conditions for the inhomogeneity and its derivatives up to high order, which are questionable in any real-world system.
Another question that is hotly debated in the ocean waves community is the nature of the transition from stable to unstable behavior, and especially how  it manifests phenomenologically in the ocean waves context. To fix ideas, consider the one-parameter family of spectra of the form $P(k)=CP_{0}(k)$ for some smooth, unimodal profile $P_{0}(k).$ It is easy to see that there exists a critical value $C_{*}$ so that if $C<C_{*}$ the spectrum is stable (in the sense of condition (4.1)), while for $C>C_{*}$ it is unstable. So the question is, how different are the solutions of (1.1) for $C=C_{*}-\epsilon$ versus $C=C_{*}+\epsilon.$ Is there a striking change in the qualitative behavior of solutions around $C=C_{*}?$
A key piece of context for this question is the timescales of interest: physically, we expect the original equation (1.1) to be valid for an $O(1)$ timescale $T,$ but not for $t\to\infty.$ In that context, instability that takes an asymptotically long time to manifest, would appear similar to stability when $t\in[0,T]$ [5]. In Section 4.2 we investigate numerically the growth in the case of weak instability.
Another aspect of this question is how large the inhomogeneity eventually becomes. In the ocean wave context, $u$ is compared to an $O(1)$ homogeneous background. Thus, even if $u$ itself changes a lot, as long as it remains dominated by the background then the phenomenology of the sea state does not change noticeably. We develop an idea of [26] and introduce the amplification factor for the problem as a way to measure the effect of the fully developed modulation instability to the overall sea state. This is discussed in some detail in Section 4.3.
A subtle issue in this discussion is the localization of the problem from full space $(x,y)\in\mathbb{R}^{2}$ to a finite domain $(x,y)\in[-\frac{L}{2},\frac{L}{2}]^{2}.$ The linear stability analysis is typically carried out on full space. In [6] the question of how to localize the Alber equation and its linear stability analysis on a finite computational domain $(x,y)\in[-\frac{L}{2},\frac{L}{2}]^{2}$ was investigated. It was found that setting
$\Gamma(x)=\frac{1}{L}\sum\limits_{n\in\mathbb{Z}}P(\frac{n}{L})e^{2\pi i\frac{n\cdot x}{L}}$ (4.3)
leads to the same stability condition as full space, when $L$ is large enough. In other words, we start from the continuous power spectrum $P(k)$ (often measured in the field in the ocean waves context), and proceed to generate an autocorrelation function $\Gamma$ according to (4.3). This in turn provides a benchmark for the size of $L$ required: as long as the effective support of $\Gamma_{\text{inf}}(x):=\mathcal{F}^{-1}_{k\to x}[P(k)]$ is well contained in $\frac{L}{2},\frac{L}{2}],$ then numerically $\Gamma_{\text{inf}}$ and the $\Gamma$ of (4.3) are interchangeable.
The original instability condition (4.1) is not straightforward to check, as it involves the existence or not of solutions in a system of two equations (real and imaginary parts of $\mathbb{H}[D_{X}P](\omega)=4\pi p/q$) in three unknowns ($X,\mathrm{Re}(\omega),\mathrm{Im}(\omega)$). A constructive way to check stability was introduced in [3]: the curve
$S_{X}(t):=\mathbb{H}[D_{X}P](t)-iD_{X}P(t),\qquad t\in\mathbb{R},$ (4.4)
is drawn on the complex plane. This is a closed curve starting and ending at 0 (and it might be self-intersecting for multimodal spectra $P$). If the winding number of $S_{X}$ around $4\pi p/q$ is nonzero, then the wavenumber $X$ is unstable. This method computes the solutions $X_{*}$ of (4.1), but it only clarifies the existence of $\omega_{*}$ – not their values. In Figure 5 this technique is demonstrated for the Gaussian spectrum (4.5) with $C=1.6.$ From this, the bandwidth of unstable wavenumbers can be computed. For the required computation of the Hilbert transform we use Weideman’s rational eigenfunction method [27]. Standalone scripts implementing this check of the stability condition can be found in https://github.com/aathanas/penrose4alber.
Finally, it is worth noting that, in the ocean waves problem, the potential $V(x,t)=u(x,x,t)$ controls the size of deviation from the homogeneous background at location $x$ and time $t.$ Thus, an appropriate way to keep track of the size of the inhomogeneity over space and time is through $|u(x,x,t)|.$
Figure 5. The curves $S_{X}(t)$ are plotted on the complex plane for various values of $X\in\mathbb{R}$ as in (4.4). In the context of Section 4.2 $p=q=1,$ hence the target point is $4\pi p/q=4\pi.$ For this particular case $C=1.6,$ the first four harmonics of the computational domain, $X=2\pi n/L,$ $n\in\{1,2,3,4\},$ are unstable. The bandwidth of unstable wavenumbers for this case is approximately $1.13.$
Figure 6. The modulus $|u(x,x,t)|$ is plotted. The inhomogeneity $u$ satisfies the fully nonlinear equation (1.1) with $p=q=1,$ $\Gamma$ consistent with equation (4.5) and $C=1.9.$ The initial condition $u_{0}$ is given in (4.6).
4.2. Numerical investigation of the bifurcation
The onset of modulation instability is not fully understood yet, but it is considered one of the possible mechanisms behind the growth of oceanic rogue waves [13]. Hence investigating its fully nonlinear behavior presents great interest both from a theoretical dynamical systems point of view, as well as an ocean waves point of view. Capturing the growth (or not) of the inhomogeneity is precisely why the discrete balance law of Proposition 2.2 is important, as it guarantees that the numerical solution qualitatively follows the exact one. Moreover, the second order in time and fourth order in space of the numerical method presented here makes tractable the large-domain and long-time computations required to conclusively investigate nonlinear Landau damping (in the stable case), and the growth of coherent structures (in the unstable case).
For clarity and compactness, here we investigate problems where all the terms are of comparable sizes, $p=q=1,$ and use a Gaussian power spectrum of
$P(k)=\frac{C^{2}}{\sigma}e^{-\pi\frac{k^{2}}{\sigma^{2}}},\qquad\sigma=0.36,\qquad C\in[0.9,2.2].$ (4.5)
The parameter $C$ controls the RMS wave amplitude in the background sea state. Moreover, this is a relatively narrow power spectrum, i.e., the situation in which modulation instability is expected to appear. We use an initial inhomogeneity of the form
$\begin{array}[]{c}f_{0}(x,y)=0.05\cdot e^{-0.06x^{2}-0.07y^{2}}(1+A_{1}\cos(0.3x)\cos(0.2y)+A_{2}x+A_{3}y),\\[4.0pt] u_{0}(x,y)=\frac{1}{2}\left(f_{0}(x,y)+\overline{f_{0}(y,x)}\right).\end{array}$ (4.6)
In order to have directly comparable results for different values of $C,$ we used
$A_{1}=0.3+0.8i,\qquad A_{2}=-0.2,\qquad A_{3}=0.1i$ (4.7)
in all the runs discussed in Section 4.2. Values drawn from a complex normal distribution for $A_{1},A_{2},A_{3}$ yield similar results, i.e., the results are not particularly sensitive to the exact values – as we will see in Section 4.3, where randomized values are used. The strongest dependence of the results on the initial data is through the overall size of the initial condition, as for smaller initial inhomogeneities it can take substantially longer for the inhomogeneity to grow in the unstable case. The initial datum of (4.6), (4.7) has $\|u_{0}\|_{L^{2}_{x,y}}\approx 0.31$ and $\|u_{0}\|_{L^{\infty}_{x,y}}\approx 0.06;$ in other words it is small but not vanishingly so. We use an initial inhomogeneity that is not symmetric nor a very simple function to highlight the robustness of the qualitative behavior of the problem. All numerical results in this Section, including Figures 6-9, are carried out with $u_{0}$ as in (4.6), (4.7), on a computational domain $(x,y)\in[-\frac{L}{2},\frac{L}{2}]^{2}$ with $L=50,$ timestep size $\tau=10^{-2}$ and space mesh size $h=7\cdot 10^{-2}.$
Figure 7. Evolution in time of the norms of the solution $u$ of the fully nonlinear problem (1.1); and $u_{lin},$ the solution of the linearized problem (4.8). Here $p=q=1,$ $\Gamma$ consistent with equation (4.5) and $C=1.9.$ The initial condition $u_{0}$ is given in (4.6).
In the case of modulation instability, the inhomogeneity first goes through a phase of exponential growth, and a localized coherent structure with values of $O(1)$ appears. Then, growth in $L^{\infty}$ sense stops, and a distinctive pattern emerges, of similar localized structures spreading outwards in space to form a lattice within a space-time cone. A typical example of the space-time lattice of coherent structures can be seen in Figure 6. This matches very closely the universal behavior of modulation instability in other wave problems [10]. The main features of this cone-supported lattice (such as the size and distance of the blobs, the maximum amplitude etc) does not seem to depend on the inhomogeneity, but only on the homogeneous background $\Gamma.$ This finding is further supported by the systematic Monte Carlo investigation reported in Figure 10. To the best of our knowledge, there are no results in finding this meta-stable pattern directly (i.e. without solving the IVP (1.1) for particular initial inhomogeneities).
When $C=1.9$ as in Figure 6, this is a relatively intense background giving rise to a relatively strong instability. Here the values of $|u|$ peak at about $6,$ clearly dominating the background of $1.9.$ The two distinct stages are visible in Figures 6 and 7: first, there is exponential growth of the inhomogeneity to a single, localized coherent structure. Once the first coherent structure is formed, there is no more growth in the $L^{\infty}$ norm, but the pattern propagates outwards forming the lattice-within-a-cone.
For purposes of comparison with the linear stability analysis of the problem, we also solve the linearized Alber equation,
$\begin{array}[]{c}i\partial_{t}u_{\text{lin}}+p(\Delta_{x}-\Delta_{y})u_{\text{lin}}+q\Bigl(u_{\text{lin}}(x,x,t)-u_{\text{lin}}(y,y,t)\Bigr)\,\Gamma(x-y)=0,\\[8.0pt] \qquad u_{\text{lin}}(x,y,0)=u_{0}(x,y).\end{array}$ (4.8)
In Figure 7 we see the evolution in time of the $L^{2}$ and $L^{\infty}$ norms for $u$ and $u_{lin}$ for $C=1.9.$ By $t=1$ or so the solutions of both the fully nonlinear and linearized equations have hit their maximum rate of growth. This picking-up-steam stage can take much longer when the instability is weaker (e.g., smaller $C$). Figure 7 shows that the linearized equation (4.8) captures very well the initial stage of exponential growth of the unstable modes. On the other hand, the linearized equation completely misses the maxing out of the $L^{\infty}$ norm and the formation of a coherent structure. In conclusion, the linear stability analysis is found to accurately predict the onset of modulation instability, as well the rate of growth in the early stage. However, it clearly misses things like the maximum amplitude $\max|u|$, the space-time cone etc. Full solution of the nonlinear equation (1.1) is required for that.
Figure 8. Evolution of the size of $u(t)$ in time for different values of $C.$ The inhomogeneity $u$ satisfies the fully nonlinear equation (1.1) with $p=q=1,$ $\Gamma$ consistent with equation (4.5) and initial condition $u_{0}$ is given in (4.6). Top: evolution in time of the $L^{2}$ norm in space. Bottom: evolution in time of the $L^{\infty}$ norm in space.
Table 6. Observed maximum inhomogeneity versus background waves $C$ for the problems of Figures 6, 9. The inhomogeneity $u$ satisfies the fully nonlinear equation (1.1) with $p=q=1,$ $\Gamma$ consistent with equation (4.5) and initial condition $u_{0}$ is given in (4.6).
$C$ $\|u\|_{L^{\infty}_{x,y,t}}$ Ratio of max inhomogeneity over background
1.9 6.6 3.47
1.6 2.4 1.5
1.5 1.5 1
1.4 0.7 0.5
1.3 0.3 0.23
In Figure 8 we can see how the qualitative behavior changes with the strength of the background $C.$ The values $C\in[1.3,1.6]$ also lead to instability, but one that grows more slowly and to a smaller maximum, as compared to $C=1.9.$ Perhaps more important than the raw amplitude of the inhomogeneity, is its size in comparison to the RMS background waves, $C$. As can be seen in Table 6, below $C=1.5$ the observed maximum amplitude of $u$ is smaller than the background $C.$ Thus, examining the case $C\in[1.3,1.9],$ we see linear instability present in all cases (and indeed the inhomogeneity grows by at least one order of magnitude from its initial size), but this does not always lead to a dominant effect when compared to the homogeneous background. In other words, the presence of modulation instability does not automatically mean that the autocorrelation changes substantially. Of course, for intense enough background, e.g. $C>1.6,$ dominant effects are clearly present.
Finally, in Figure 9 we see a typical example of nonlinear Landau damping for the equation (1.1). There is no modulation instability present, only dispersion. In this case, the linearized equation (4.8) provides an excellent approximation to the fully nonlinear one.
Figure 9. The modulus $|u(x,x,t)|$ is plotted. The inhomogeneity $u$ satisfies the fully nonlinear equation (1.1) with $\Gamma$ consistent with equation (4.5) and $C=0.9.$ The initial condition $u_{0}$ is given in (4.6). Observe that the solution does not grow significantly from its initial magnitude.
4.3. Monte Carlo investigation of the amplification factors
If we go back to the derivation of equation (1.1), the total second-moment of the underlying sea state is given by
$\rho(x,y,t)=u(x,y,t)+\Gamma(x-y).$ (4.9)
In that setting, the quasi-homogeneity assumption is reflected in the scaling $u(x,y,0)=o(\Gamma).$ Thus, the growth of the inhomogeneity $u$ can lead to significant localized events only if $u(x,y,t)$ becomes comparable to or larger than $\Gamma.$ Following [26, 23], we keep track of the quantity
$\mathcal{A}:=\frac{\max\limits_{t,x,y}|u(x,y,t)+\Gamma(x-y)|}{\max\limits_{x,y}|\Gamma(x-y)|}=\frac{\max\limits_{t,x,y}|u(x,y,t)+\Gamma(x-y)|}{\Gamma(0)}$ (4.10)
which we call the amplification factor for the problem, as it controls how much larger localized extreme events can be as compared to the homogeneous background. This quantity is referred to as $\tilde{\rho}(0,0,\tilde{\tau})/\tilde{\rho}_{h}(0)$” in [26, 23], and in [23] the rogue wave condition is reported to be equivalent to
$\mathcal{A}>2.84.$ (4.11)
It is crucial to understand how and when large amplification factors are possible. Moreover, in the context of ocean waves, amplification factors close to $1$ are essentially equivalent to stability – even if technically modulation instability is present.
Table 6 suggests that, in the context of the spectrum (4.5), amplification factors of at least $\mathcal{A}\approx 1.5$ are to be expected when $C\approx 1.4.$ As a starting point, we want to see how sensitive this threshold is with regard to the exact initial condition. After all, the exact inhomogeneity in the field is not knowable – Table 6 makes sense to the extent that it is representative “for every reasonable inhomogeneity”.
Therefore, we perform a Monte Carlo investigation, and compute amplification factors for different values of $C$ in (4.5), and different values of the coefficients $A_{j}$ in equation (4.6). More specifically, we set
$C\sim\mathcal{U}([0.9,2.2]),\qquad A_{j}\sim\mathcal{N}(0,\frac{1}{9})+i\mathcal{N}(0,\frac{1}{9}).$ (4.12)
The resulting problem of the form (1.1) is then solved for $t\in[0,14],$ on $x,y\in[-L,L]^{2}$ with $L=20,$ and discretization parameters $\tau=10^{-2},$ $h=7\cdot 10^{-2}.$ This is the same resolution as in Section 4.2, on a slightly smaller time and space window, for reasons of computation time. Still, there is sufficient time and space for the modulation instability to fully develop. The resulting amplification factors, plotted against the background strength $C,$ can be seen in Figure 10.
The numerical simulations for Figure 10 were performed on the standard nodes of Hypatia HPC cluster at the University of St Andrews (each standard node contains two 84-core AMD EPYC 9634 CPUs, 384GB of memory and a local 1.7TB scratch disk). We used MATLAB R2025a with SLURM for job scheduling. We executed $150$ independent realizations of the PDE initial value problem solver in batches of 10 parallel processes, each allocated 6 CPU cores to maximize BLAS efficiency.
It is clearly seen that the amplification factor is not sensitive to the exact details of the initial inhomogeneity; moreover, a robust dependence on $C$ clearly emerges. To the best of our knowledge, this dependence of the maxima of the fully developed modulation instability on the background strength and not on the initial condition, was not known before.
Figure 10. Amplification factors $\mathcal{A}$ as defined in (4.10) for the problem specified in (4.5), (4.6) with parameters randomized according to (4.12). It must be noted that modulation instability onset is at $C\approx 0.9,$ but amplification factors substantially larger than 11 first appear for $C\approx 1.35$ and grow linearly past that point.
In light of Figure 10, there are several key values for the strength of the background wavefield, $C.$ First, at $C\approx 1,$ linear instability kicks in. However, at those borderline unstable cases, the inhomogeneity plateaus to values that are small compared to the background. Thus, phenomenologically, weak instability presents similarly to stability. For $C>1.35,$ we have amplification factors that are meaningfully larger than $1,$ thus the effect of the instability is noticeable. These maxima are recurrent in time over the origin $x=0$ and also would be expected to be recurrent in space after some time (cf. e.g. Figure 6). Random fluctuations of sea states with these non-homogeneous autocorrelations could conceivably generate rogue waves. Finally, for $C>1.9,$ we observe that $\mathcal{A}>2.84$ is expected with virtual certainty. This would mean regularly recurrent rogue waves, and appears almost at double the intensity of onset of linear instability. It is possible that this latter case is too strong, and does not actually describe a realistic situation in the ocean.
It is worth noting that around $C\approx 1.35$ the bandwidth of unstable wavenumbers becomes approximately $1,$ i.e. comparable to the peak wavenumber in this problem (since $p=q=1$). This, i.e. the unstable bandwidth becoming comparable to $k_{0},$ may be a condition for the second, phenomenological bifurcation to appear.
5. Discussion and Outlook
5.1. The problem and the new numerical method
The von Neumann equation with power nonlinearity is widely used as a model for the second moments of stochastic ocean wave fields. It is known to exhibit a bifurcation between a stable regime akin to Landau damping, and a modulationally unstable regime. While a large number of works investigates this bifurcation [1, 5, 3, 14, 25, 26], and its consequences for the generation of rogue waves [13, 23], very little exists in terms of full numerical solutions of the equation.
We introduce a relaxation-Crank-Nicolson scheme for the von Neumann equation with power nonlinearity. The scheme is second-order in time, and requires the solution of a linear system in each timestep, i.e., it is explicit in the nonlinearity. This is achieved by the introduction of an auxiliary variable on a staggered time-grid. The scheme is implemented with fourth-order finite differences in space. It inherits at the discrete level the problem’s $L^{2}$ balance law (Proposition 2.2), making it particularly well suited to investigate the possible growth of the solution.
In Section 3 the scheme is validated with a soliton-derived exact solution. It is found that the initialization of the auxiliary variable is important so that the scheme achieves the expected order in both space and time, for both the main and auxiliary variables. Moreover, it is found that the main variable is to a surprising extent not affected by the error in the auxiliary variable, for example exhibiting the correct order of convergence even if the auxiliary variable does not (Tables 3 and 4). This seems to indicate some kind of averaging, and understanding it fully could be vital in fully analyzing this kind of scheme. Finally, in Figure 4 the bulk of the error is seen to be due to error in phase velocity. This is to be expected from the von Neumann stability analysis for the linear part of the equation (where our method becomes the Padé(2,2) method) – and the numerical validation shows that it is preserved in the nonlinear setting too.
A standalone version of the solver, including implementations of many of the features discussed here (EOC, constraint error, invariants, growth of inhomogeneity, Monte Carlo investigation of the amplification factor) can be found in https://github.com/aathanas/rcn_Alber_MATLAB.
5.2. New findings on the bifurcation
In Section 4 the bifurcation between Landau damping and modulation instability is discussed and numerically investigated. If the homogeneous background is smaller than the linear instability threshold ($C\approx 0.9$ in this case), there is no modulation instability at all. Instead, dispersion dominates, and the linearized Alber equation (4.8) stays very close to the fully nonlinear dynamics of (1.1).
For somewhat more intense backgrounds we get into in the borderline unstable cases. he prediction that, for weak instabilities, the inhomogeneity grows very slowly [5] is confirmed. A novel finding is that in the borderline unstable cases the inhomogeneity maxes out to amplitudes that are small compared to the homogeneous background (even if it grows dramatically compared to its initial condition). To explore this question in depth, we introduce the concept of amplification factor, i.e. how much larger the total autocorrelation becomes from the homogeneous background. This is the starting point for the Monte Carlo investigation described below.
For somewhat stronger instabilities, all the standard signs of modulation instability appear quickly and clearly. The fully non-linear space-time cone [10] is recovered, cf. Figure 6. This includes recurrence of maxima in time over $x=0$ reported in [26, 23]. Furthermore, here we can see the spreading out in space of the same pattern (under periodic boundary conditions).
Thus, the linear stability analysis [1, 3] is found to accurately capture the onset and initial growth of modulation instability. On the other hand, key artifacts are recovered by numerical solution of the fully nonlinear problem (such as the formation of a coherent structure, or the maximum amplitude of the inhomogeneity) that are inherently inaccessible by the linearized analysis.
In Section 4.3, a Monte Carlo study including stable and unstable spectra is described in detail. The point is to understand the behavior of the total field $u+\Gamma,$ quantified by the amplification factor, defined in (4.10). (In particular, large growth of $u$ in time is still phenomenologically negligible if it stays small with regard to $\Gamma$.) There it is found that the amplification factor does not depend on the initial inhomogeneity, but only on the homogeneous background. Since inhomogeneities are not typically known or measured in the field, this feature is vital for the applicability of the Alber equation in realistic ocean waves. Moreover, looking at the amplification factors, the key change in behavior is not at the onset of modulation instabilty at $C\approx 0.99,$ but when the instability is strong enough to produce noticeable inhomogeneous events at $C\approx 1.35,$ cf. Figure 10.
Thus, a new bifurcation is discovered, around $C\approx 1.35.$ For this case study, this represents a problem with $35\%$ greater intensity than the one required for the onset of modulation instability $C\approx 0.99.$ Keeping in mind e.g. Figure 5 of [3], this is plausible for the ocean, and can be a contributing factor to rogue waves. This finding reframes the question of what role the modulation instability plays in real-world ocean waves: merely the onset of modulation instability is not expected to produce significant effects. Instead, significant effects are only expected for strong enough modulation instability.
Appendix A Derivation of invariants
The following invariants for (1.1) were introduced in [23]. We include them here, along with their derivation in our notation, for completeness.
Lemma A.1.
Let $u$ satisfy equation (1.1) with Hermitian initial data, and denote
$I_{1}[u]$ $:=\int\limits_{x}u(x,y,t)\Big|_{y=x}dx,$ (A.1)
$I_{2}[u]$ $:=\int\limits_{x}(\partial_{x}-\partial_{y})u(x,y,t)\Big|_{y=x}dx,$ (A.2)
$I_{3}[u]$ $:=\left(\frac{q}{p}\int\limits_{x}u^{2}(x,y,t)\Big|_{y=x}dx+\int\limits_{x}(\partial_{x}-\partial_{y})^{2}u(x,y,t)\Big|_{y=x}dx\right).$ (A.3)
Then,
$\frac{d}{dt}I_{j}[u]=0\qquad\mbox{ for }j\in\{1,2,3\}.$ (A.4)
Remark: The function $u(x,x,t)$ is real-valued by the symmetry of the problem, but not non-negative. Thus the conservation of $I_{1}[u]$ is compatible with potentially large values of $u,$ but there will have to be positive and negative values so that the total integral is preserved.
Proof: To prove this lemma, it is helpful to use the change of variables
$\left\{\begin{array}[]{c}X=\frac{x+y}{2}\\ r=x-y\end{array}\right\}\iff\left\{\begin{array}[]{c}x=X+\frac{r}{2}\\ y=X-\frac{r}{2}\end{array}\right\}.$ (A.5)
Originally $x,y$ are two position variables, and $X$ has the physical meaning of “mean position” while $r$ of “lag” between the two positions. For any $j=1,2,\dots,$ observe that $\partial_{x}^{j}=(\partial_{X}+\frac{1}{2}\partial_{r})^{j},$ $\partial_{x}^{j}=(\partial_{X}-\frac{1}{2}\partial_{r})^{j}.$
Now denote
$f=f(X,r,t)=u(x,y,t),\qquad n(X,t):=f(X,0,t).$ (A.6)
So if $u(x,y,t)$ satisfies equation (1.1), then $f$ satisfies
$if_{t}-p\partial_{X}\cdot\partial_{r}f-q\Bigl(n(X+\frac{r}{2})-n(X-\frac{r}{2})\Bigr)\,\Bigl(\Gamma(r)+f\Bigr)=0.$ (A.7)
In this formalism, equation (A.4) is elaborated to
$\frac{d}{dt}\int\limits_{X}n(X,t)dX=0,$ (A.8)
$\frac{d}{dt}\int\limits_{X}\partial_{r}f(X,r,t)\Big|_{r=0}dX=0,$ (A.9)
$\frac{d}{dt}\left(\frac{q}{p}\int\limits_{X}n^{2}(X,t)dX+\int\limits_{X}\partial_{r}^{2}f(X,r,t)\Big|_{r=0}dX\right)=0$ (A.10)
Proof of equation (A.8): setting $r=0$ in equation (A.6) and integrating in $X$ leads to
$\int\limits_{X}\Bigl(if_{t}(X,0,t)-p\partial_{X}\partial_{r}f(X,0,t)\Bigr)dX=0\implies\frac{d}{dt}\int\limits_{X}f(X,0,t)dX=0$
by performing integration by parts in $X.$
Proof of equation (A.9): differentiating equation (A.7) with respect to $r_{j}$ yields
$\begin{array}[]{c}i\partial_{t}f_{r_{j}}-p\partial_{X}\cdot\partial_{r}f_{r_{j}}-\frac{q}{2}\Bigl(n_{X_{j}}(X+\frac{r}{2},t)+n_{X_{j}}(X-\frac{r}{2},t)\Bigr)\,\Bigl(\Gamma(r)+f\Bigr)\hfill\\ \hfill-q\Bigl(n(X+\frac{r}{2})-n(X-\frac{r}{2})\Bigr)\,\Bigl(\partial_{r_{j}}\Gamma(r)+f_{r_{j}}\Bigr)=0.\end{array}$ (A.11)
Now setting $r=0$ this becomes
$i\,\partial_{t}f_{r_{j}}(X,0,t)-p\,\partial_{X}\!\cdot\!\partial_{r}f_{r_{j}}(X,0,t)-q\bigl(\Gamma(0)+n(X,t)\bigr)\,n_{X_{j}}(X,t)=0.$
Integrating over $X$ gives
[…]
Equation (A.9) follows by integration by parts in the two last terms.
Proof of equation (A.10): by differentiating twice in $r,$ equation (A.7) becomes
$\begin{array}[]{r}i\partial_{t}\Delta_{r}f-p\partial_{X}\cdot\partial_{r}\Delta_{r}f-\frac{q}{4}\Bigl(\Delta_{X}n(X+\frac{r}{2})-\Delta_{X}n(X-\frac{r}{2})\Bigr)\Bigl(\Gamma(r)+f(X,r,t)\bigr)\\ -q\nabla_{X}\Bigl(n(X+\frac{X}{2})+n(X-\frac{n}{2})\Bigr)\cdot\nabla_{r}\Bigl(\Gamma(r)+f(X,r,t)\Bigr)\\ -q\Bigl(n(X+\frac{r}{2})-n(X-\frac{r}{2})\Bigr)\Delta_{r}\Bigl(\Gamma(r)+f(X,r,t)\Bigr)=0.\end{array}$
By setting $r=0$ we get
$\begin{array}[]{r}i\partial_{t}\Delta_{r}f(X,0,t)-p\partial_{X}\cdot\partial_{r}\Delta_{r}f(X,0,t)\hfill\\ -q\nabla_{X}\Bigl(n(X+\frac{X}{2})+n(X-\frac{n}{2})\Bigr)\cdot\nabla_{r}\Bigl(\Gamma(r)+f(X,0,t)\Bigr)=0;\end{array}$
by integrating in $X$ and some obvious integrations by parts
$\begin{array}[]{r}i\frac{d}{dt}\int\limits_{X}\Delta_{r}f(X,0,t)dX-2q\int\limits_{X}\nabla_{X}n(X,t)\cdot\nabla_{r}f(X,0,t)dX=0.\end{array}$ (A.12)
Moreover, by multiplying equation (A.7) with $f,$ setting $r=0$ and integrating in $X$ we get
$\begin{array}[]{c}\int\limits_{X}f(X,0,t)\Bigl(if_{t}(X,0,t)-p\partial_{X}\cdot\partial_{r}f(X,0,t)\Bigr)dX=0\implies\hfill\\ \hfill\implies-\frac{i}{p}\frac{1}{2}\frac{d}{dt}\int\limits_{X}n^{2}(X)dX=\int\limits_{X}\nabla_{X}n(X,t)\cdot\nabla_{r}f(X,0,t)dX\end{array}$ (A.13)
By combining equations (A.12) and (A.13), relation (A.10) follows. ∎